OpenStudy (anonymous):

Convert 3x - 4y = 2 into a slope-intercept form.

OpenStudy (anonymous):

@sauravshakya Help ?

OpenStudy (anonymous):

Take it to the from y=mx+c

OpenStudy (anonymous):

yup bu idk how to. :( cuz instead of plus its a subtract.

OpenStudy (anonymous):

3x - 4y = 2 -4y=-3x+2 4y=3x-2 y=3/4 x -2/4 y=3/4x -1/2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yup.